If an object is propelled upward from ground level with an initial velocity of 224 feet per second, its height h in feet t seconds later is given by the equation h = 224t - 16t2. After how many seconds does the object hit the ground?
Answer by stanbon(75887)

If an object is propelled upward from ground level with an initial velocity of 224 feet per second, its height h in feet t seconds later is given by the equation h = 224t - 16t2. After how many seconds does the object hit the ground?
---
It hits the ground when its height is zero.
--------
Solve -16t^2 + 222t = 0
Factor:
-2t(8t- 111) = 0
t = 0 means the object was on the ground at time zero.
t = 111/8 = 13.875 sec. is the time when the object is on the ground again.
=============================
Cheers,
Stan H.